OpenStudy (anonymous):

Simplify the rational expression state any excluded values. A. 2 x-8/x-4 B. 4x-8/4x+20

OpenStudy (anonymous):

\[\frac{ 2x-8 }{ x-4 }\]

OpenStudy (anonymous):

b. \[\frac{ 4x-8 }{ 4x+20 }\]

OpenStudy (anonymous):

Hey @EmeraldSunset do you know how to do these?

OpenStudy (anonymous):

@dumbcow @phi i really need help!

OpenStudy (phi):

in 2x -8 what number can you divide into each term ?

OpenStudy (anonymous):

by 2

OpenStudy (anonymous):

so you end up with\[x-4\]

OpenStudy (phi):

yes, you can divide 2x-8 by 2 you can write it as 2(x-4) (notice that if you "distribute" the 2, you get 2x-8, which is what you started with) now use the idea that anything divided by itself is 1 (except zero divided by zero) \[ \frac{2(x-4)}{x-4} = 2 \cdot \frac{x-4}{x-4}= 2\cdot 1 = 2\] there is one complication: 0/0 is not defined. If x=4 we would have 4-4/(4-4)= 0/0 and that is not allowed. To handle that, we say the answer is 2, with the restriction x≠ 4